Vaughn Corporation purchased a new machine for its assembly process on August 1, 2020. The cost of this machine was $153,300. The company estimated that the machine would have a salvage value of $18,300 at the end of its service life. Its life is estimated at 5 years, and its working hours are estimated at 22,500 hours. Year-end is December 31.

Compute the depreciation expense under the following methods. Each of the following should be considered unrelated. (Round depreciation rate per hour to 2 decimal places, e.g. 5.35 for computational purposes. Round your answers to 0 decimal places, e.g. 45,892.)

(a)

Straight-line depreciation for 2020

$enter a dollar amount

Entry field with incorrect answer

(b)

Activity method for 2020, assuming that machine usage was 800 hours

$enter a dollar amount

Entry field with correct answer

(c)

Sum-of-the-years'-digits for 2021

$enter a dollar amount

Entry field with incorrect answer

(d)

Double-declining-balance for 2021
Answer #1

a.

.Straight line method = (Cost - Salvage value) / years

= ($153,300 - $18,300) / 5 years

= $27,000

Depreciation expense for 5 months (From August to Dec. 2020) = $27,000 * (5/12)

= $11,250

b.

Activity method = (2020 working hours / Total working hours) * (Cost - salvage value)

= (800 / 22,500) * ($153,300 - $18,300)

= $4,800

c.

Sum of years digit method

5 + 4 + 3 + 2 + 1 = 15

Depreciation for first 7 months (From January 2021 To August 2021) = (Cost - Salvage value) * (5/15) * (7/12)

= ($153,300 - $18,300) * (5/15) * (7/12)

= $26,250

Depreciation for next 5 months (From August 2021 to December 2021) = (Cost - Salvage value) * (4/15) * (5/12)

= ($153,300 - $18,300) * (4/15) * (5/12)

= $15,000

Total Depreciation = $26,250 + $15,000

= $41,250

d.

Rate of depreciation = (100/5) * 2

= 40%

Depreciation for the First 5 months From August 2020 to December 2020

= cost * (5/12) * 40%

= $153,300 * (5/12) * 40%

= $25,550

Depreciation for the next 7 months From January 2021 to Jul 2021

= cost * (7/12) * 40%

= $153,300 * (7/12) *40%

= $35,770

Depreciation from August 2021 to December 2021

= ($153,300 - $25,550 - $35,770) * (5/12) * 40%

= $91,980 * (5/12) * 40%

= $15,330

Total depreciation 2021 = $35,770 + $15,330

= $51,100